# Leveraging Offline Data in Linear Latent Bandits

### v1

@@ 8 Discussion, Limitations and Further Work
In this paper, we have addressed the problem of leveraging offline data to accelerate online learning in linear latent bandits. Our work has a few limitations. First, while ProBALL-UCB is practical and computationally efficient, it has a slightly weaker worse-case guarantee than the less practical LOCAL-UCB. Second, when the size $N$ of the offline dataset is small, it can be hard to tune the regularization $\mu$. We use a pseudoinverse-based version of SOLD in such a case, implemented in our code. Third, the offline uncertainty sets computed using $\Delta_{\operatorname{off}}$ can be overly conservative, and a discount hyperparameter $\tau$ must be fine-tuned online.
Despite these limitations, our work enjoys strong theoretical guarantees and convincing empirical performance. We hope that this method opens the door for developing other efficient and scalable algorithms for sequential decision-making with continuous latent states. One can use ideas presented in this paper to design similar algorithms for MDPs, linear MDPs as well as under general function approximation for both MDPs and bandits. It is also possible that our regret bounds can be sharpened using feature-space partition ideas from hybrid RL (Li et al., 2023; Tan and Xu, 2024). Finally, it will be of interest to obtain minimax guarantees for leveraging offline data to speed up online learning in latent and linear latent contextual bandits.
@@ A.2 Coherence with Equality is Necessary
Consider an SDP with action set $\mathcal{A}=\{0,1\}$ equipped with a random variable $\theta\in\{0,1\}$ distributed as $Ber(1/2)$. If the first action is 00, then all rewards are given by $\theta$. If the first action is $1$, then the first reward $X_{1}=\theta$ while all future rewards are $1-\theta$.
@@ A.3 Contexts, Latent State and Behavior Policy cannot be Dependent
  • Latent contextual bandit $1$: Let $\bm{\theta}$ take values $e_{1}+e_{2}$ and $-e_{1}-e_{2}$ with probability $1/2$ each.
@@ A.3.1 Contexts cannot be dependent on each other
We consider two actions, so $\mathcal{A}=\{0,1\}$ We have two contexts $x,y$ so that $\phi(x,0)=e_{1}$, $\phi(x,1)=2e_{1}$ while $\phi(y,0)=e_{2}$, $\phi(y,1)=2e_{2}$. We design them to be dependent so that any trajectory either only sees $x$ or only sees $y$. However, $x$ and $y$ are both seen with probability $1/2$. Pick $\pi_{b}$ so that it takes each action with probability $1/2$. Now note that the mean reward of $x,0$ in either latent bandit takes values ±$\pm 1$ with probability $1/2$. So, the offline data distributions are also indistinguishable and every context-action pair is seen with probability at least $1/4$.
@@ Appendix B A de Finetti Theorem for decision processes
B.1 Discussion on Liu et al. [2023]’s smaller class of generalized bandits
We prove the de Finetti theorem for a very general formulation of decision processes. However, past work [Liu et al., 2023] has studied a simpler generalization of bandits, namely a stochastic process valued in $\mathbb{R}^{\mathcal{A}}$. A sample point in this space is a sequence of functions in $\mathcal{A}\to\mathbb{R}$, which rules out the possibility of adaptivity. In contrast, a sample point in our space is a function of sequences, which subsumes all sample points of their space, but allows for adaptivity.
Liu et al. [2023] are able to use the original de Finetti theorem on their random process directly, but work with a much more restrictive kind of decision process. We show that even when considering much more general stateless decision processes, latent bandits are the "right" objects produced by a de Finetti theorem for stateless decision processes.
@@ B.2 Proof of the De Finetti Theorem for Stateless Decision Processes
We recall our de Finetti theorem for stateless decision processes here. See 1
@@ B.3 Constructing $F$
The trick: Consider the infinite sequence $\tau_{\infty}:=(a_{h},a_{h},\dots)$. By coherence, $Y_{\tau,h}=Y_{\tau_{\infty},h}$. Now $\tau_{\infty}$ is a sequence where exchanging any finite set of rewards preserves the reward distribution, since all actions are identical. Since $\mathbb{R}$ is locally compact, we can apply the usual de Finetti representation theorem (Theorem 12.26 in Klenke [2008]) to conclude that:
@@ B.4 Establishing that $\mathcal{L}[Y_{\tau,h}\mid F]=F(a_{h})$ almost surely for all $h$
[…]
We will show our claim using the definition and a.s. uniqueness of the regular conditional distribution in our case. Consider any $F$-measurable set $A\subset\Omega$ and Borel-measurable set $B\subset\mathbb{R}$, $B\in\mathcal{B}$. Denote by $\kappa_{F}:=\mathcal{L}[Y_{\tau,h}\mid F]$ and by $\kappa_{a_{h}}:=\mathcal{L}[Y_{\tau,h}\mid F(a_{h})]=F(a)$. Note that by the coherence and exchangeability in section B.3,
[…]
@@
[…]
@@
where the last equality holds by the dominated convergence theorem, if the limit exists. To establish that the limit exists and compute it, we apply the usual de Finetti theorem – specificaly point $(i)$ of remark 12.27 in Klenke [2008] with $f$ set to the identity map. This gives us that $\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\mathds{1}_{B}(Y_{\tau_{\infty},j})=\mathbb{E}[\mathds{1}_{B}(Y_{\tau_{\infty},h})\mid\Xi_{a_{h}}]=\Xi_{a_{h}}(B)$, where $\Xi_{a_{h}}(B)$ is the random variable $\omega\mapsto\Xi_{a_{h}}(\omega)(B)$. This in turn satisfies $\Xi_{a_{h}}(\omega)(B)=F(a_{h})(\omega)(B)=\kappa_{a_{h}}(\omega,B)$. This establishes that for any $F$-measurable set $A$ and Borel set $B$,
[…]
@@
[…]
@@ B.5 Establishing conditional independence of rewards
We will show this by induction on $H$. This clearly holds for $H=1$ by section B.3 above. Now assume that this holds for $H=l-1$.
@@
Showing that the empirical average is the conditional expectation: Again, consider a different infinite trajectory $\tau_{l}=(a_{l},a_{l},\dots)$. Again, by coherence, $Y_{\tau_{l},l+j}=Y_{\tau^{\prime},l+j}$ for all $j\geq 0$. From the usual de Finetti theorem, specifically point (i) of remark 12.27 in Klenke [2008], we have that $\frac{1}{m}\sum_{j=1}^{m}f_{l}(Y_{\tau_{l},j})\to\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]$. We can then observe that by general properties of convergence, the following holds.
@@
Since both $U$ and $\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]$ are now $F$ measurable, we can apply the induction hypothesis to $\tau^{\prime}$ truncated at $l-1$ and conclude that
@@ B.6 A de Finetti theorem for TACDPs
A latent contextual bandit is a stateless decision process equipped with a random measure-valued function $F:\Omega\to((\mathcal{X}\times\mathcal{A})\to\mathcal{P}(\mathbb{R}))$ so that for any $H$ and context-action sequence $((x_{1},a_{1}),\dots(x_{H},a_{H}))$, the rewards $(Y_{1},\dots Y_{H}):=\mathcal{F}_{H}((x_{1},a_{1}),\dots(x_{H},a_{H}))$ are independent conditioned on $F$. Moreover, the conditional distribution $\mathcal{L}[Y_{h}\mid F]=F((x_{h},a_{h}))$ for all $h\leq H$ [^6] We abuse notation twice here. First, we write $F((x_{h},a_{h})):=(\omega\mapsto F(\omega)((x_{h},a_{h})))$. Second, as the regular conditional distribution $\mathcal{L}[Y_{h}\mid F]$ is a kernel that maps from $\Omega\times\mathcal{B}\to\mathbb{R}$, we view $F((x_{h},a_{h}))$ as its curried map $(\omega,B)\mapsto F((x_{h},a_{h}))(\omega)(B)$. A discussion of issues like measurability and well-definedness is in Appendix B.2.
@@ Appendix C Proofs for SOLD
With probability at least $21-\delta/$ 2, we have that $\|\overline{\textbf{M}}_{N}-\mathbb{E}[\textbf{M}_{1}]\|_{2}\leq\Delta_{M}$ with
@@
Now since $\|\textbf{M}_{n}^{2}\|_{2}\leq\|\textbf{M}_{n}\|_{2}^{2}\leq R^{4}\left(2+\frac{H}{4\mu}\right)^{2}$, we have that $\|\textbf{M}_{n}^{2}-\mathbb{E}[\textbf{M}_{1}^{2}]\|_{2}\leq 2R^{4}\left(2+\frac{H}{4\mu}\right)^{2}$ by the matrix Hoeffding bound, we have that with probability $21-\delta/$ 2,
@@
Further, by the matrix Bernstein inequality, we have that with probability $21-\delta/$ 2,
@@
Combining the two results and using a union bound, we get that with probability $21-\delta/$ 2
@@
With probability $41-\delta/$ 4, for $i=1,2$, we have that $\|\overline{\textbf{D}}_{N,i}-\mathbb{E}[\textbf{D}_{n,i}]\|_{2}=\left\|\frac{1}{N}\sum_{n=1}^{N}\textbf{D}_{n,i}-\mathbb{E}[\textbf{D}_{n,i}]\right\|_{2}\leq\Delta_{D}$ with
@@
Now note that by inequality (1.1) from Wei et al. [2005], we have that
@@
First note that $\mathbb{E}[\textbf{D}_{N,1}]^{-1}\mathbb{E}[\textbf{M}_{N,1}]\mathbb{E}[\textbf{D}_{N,2}]^{-1}=\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. Also recall that $\hat{\textbf{U}}$ is given by the top-$d_{K}$ eigenvectors for $\mathbb{E}[\textbf{D}_{N,1}^{-1}\textbf{M}_{N,1}\textbf{D}_{N,2}^{-1}]$. This means that there is a $d_{K}\times d_{K}$ unitary matrix W such that $\textbf{U}_{\star}\textbf{W}$ forms the eigenvectors for $\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. This means that by the Davis-Kahan theorem for statisticians [Yu et al., 2014] and Lemma 5, we have that with probability $1-\delta$
@@
Bounding $\Delta_{\operatorname{off}}$: Now note that for large enough $N$, $\|\mathbb{E}[\textbf{D}_{n,1}]^{-1}\|_{2}\Delta_{D}=\sqrt{\frac{8\log(d_{A}/\delta)}{\lambda_{A}N}}\leq\frac{1}{2}$. In particular, by applying inequality (1.1) from Wei et al. [2005], we have that $B_{D}\leq 2\|\mathbb{E}[\textbf{D}_{n,1}]^{-1}\|_{2}=\frac{2}{\lambda_{A}}$. This already gives us the much simpler expression
@@ Appendix D Proofs for LOCAL-UCB
  • Note that with probability $31-\delta/$ 3, $\textbf{U}_{\star}$ satisfies $\|\hat{\textbf{U}}^{\top}\textbf{U}_{\star}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}$.
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\tilde{\textbf{U}}_{t}\tilde{\textbf{U}}_{t}^{\top}\tilde{\bm{\beta}}_{t}-\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\bm{\beta}_{\star}$
@@
=$\stackrel{{\scriptstyle(ii)}}{{=}}RT\Delta_{\operatorname{off}}+\sum_{t=1}^{T}\alpha_{2,t}\|\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\|_{\textbf{V}_{1,t}^{-1}}$
@@
=$\stackrel{{\scriptstyle(iv)}}{{=}}O\left(Rd_{K}\sqrt{T\log\left(T/\delta\right)}+\frac{R\lambda_{D}^{3}}{\lambda_{\min}}\sqrt{\frac{T^{2}d_{K}d_{A}\log(d_{A}/\delta)}{N}}\right)$

### v2

@@ 9 Discussion, Limitations and Further Work
In this paper, we have addressed the problem of leveraging offline data to accelerate online learning in linear latent bandits. Our work has a few limitations. First, while ProBALL-UCB is practical and computationally efficient, it has a slightly weaker worse-case guarantee than LOCAL-UCB. Second, when the data is noisy, it can be hard to tune the regularization $\mu$. We use a pseudoinverse-based version of SOLD in such a case (Appendix G), implemented in our code. Third, the offline uncertainty sets computed using $\Delta_{\operatorname{off}}$ can be overly conservative, and a discount hyperparameter $\tau$ must be fine-tuned online.
Despite these limitations, our work enjoys strong theoretical guarantees and convincing empirical performance. We hope that this method opens the door for developing other efficient and scalable algorithms for sequential decision-making with continuous latent states. One can use ideas presented in this paper to design similar algorithms for MDPs, linear MDPs, and RL or bandits with general function approximation.
@@ A.2 Coherence with Equality is Necessary
Consider an SDP with action set $\mathcal{A}=\{0,1\}$ equipped with a random variable $\theta\in\{0,1\}$ distributed as $Ber(1/2)$. If the first action is 0, then all rewards are given by $\theta$. If the first action is $1$, then the first reward $X_{1}=\theta$ while all future rewards are $1-\theta$.
@@ A.3 Contexts, Latent State and Behavior Policy cannot be Dependent
  • Latent contextual bandit $1$: Let $\bm{\theta}$ take values $e_{1}+e_{2}$ and $e_{1}-e_{2}$ with probability $1/2$ each.
@@ A.3.1 Contexts cannot be dependent on each other
We consider two actions, so $\mathcal{A}=\{0,1\}$ We have two contexts $x,y$ so that $\phi(x,0)=e_{1}$, $\phi(x,1)=2e_{1}$ while $\phi(y,0)=e_{2}$, $\phi(y,1)=2e_{2}$. We design them to be dependent so that any trajectory either only sees $x$ or only sees $y$. However, $x$ and $y$ are both seen with probability $1/2$. Pick $\pi_{b}$ so that it takes each action with probability $1/2$. Now note that the mean reward of $x,0$ in either latent bandit takes values $\pm 1$ with probability $1/2$. So, the offline data distributions are also indistinguishable and every context-action pair is seen with probability at least $1/4$.
@@ Appendix B A de Finetti Theorem for decision processes
B.1 Discussion on Liu et al., (2023)’s smaller class of generalized bandits
We prove the de Finetti theorem for a very general formulation of decision processes. However, past work (Liu et al.,, 2023) has studied a simpler generalization of bandits, namely a stochastic process valued in $\mathbb{R}^{\mathcal{A}}$. A sample point in this space is a sequence of functions in $\mathcal{A}\to\mathbb{R}$, which rules out the possibility of adaptivity. In contrast, a sample point in our space is a function of sequences, which subsumes all sample points of their space, but allows for adaptivity.
Liu et al., (2023) are able to use the original de Finetti theorem on their random process directly, but work with a much more restrictive kind of decision process. We show that even when considering much more general stateless decision processes, latent bandits are the "right" objects produced by a de Finetti theorem for stateless decision processes.
@@ B.2 Proof of the De Finetti Theorem for Stateless Decision Processes
We recall our de Finetti theorem for stateless decision processes here. See 5
@@ B.3 Constructing $F$
The trick: Consider the infinite sequence $\tau_{\infty}:=(a_{h},a_{h},\dots)$. By coherence, $Y_{\tau,h}=Y_{\tau_{\infty},h}$. Now $\tau_{\infty}$ is a sequence where exchanging any finite set of rewards preserves the reward distribution, since all actions are identical. Since $\mathbb{R}$ is locally compact, we can apply the usual de Finetti representation theorem (Theorem 12.26 in Klenke, (2008)) to conclude that:
@@ B.4 Establishing that $\mathcal{L}[Y_{\tau,h}\mid F]=F(a_{h})$ almost surely for all $h$
    $\mathbb{E}[\mathds{1}_{B}(Y)\mathds{1}_{A}]=\int\mathds{1}_{B}(Y(\omega))\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)=\int\kappa_{G,X}(\omega,B)\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)$
We will show our claim using the definition and a.s. uniqueness of the regular conditional distribution in our case. Consider any $F$-measurable set $A\subset\Omega$ and Borel-measurable set $\subset\mathbb{R}, BB\in\mathcal{B}$. Denote by $\kappa_{F}:=\mathcal{L}[Y_{\tau,h}\mid F]$ and by $\kappa_{a_{h}}:=\mathcal{L}[Y_{\tau,h}\mid F(a_{h})]=F(a)$. Note that by the coherence and exchangeability in section B.3,
$\int\kappa_{F}(\omega,B)\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)=\mathbb{E}[\mathds{1}_{B}(Y_{\tau,h})\mathds{1}_{A}]=\mathbb{E}[\mathds{1}_{B}(Y_{\tau_{\infty},h})\mathds{1}_{A}]=\mathbb{E}[\mathds{1}_{B}(Y_{\tau_{\infty},j})\mathds{1}_{A}]$
@@
$\int\kappa_{F}(\omega,B)\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)=\mathbb{E}[\mathds{1}_{B}(Y_{\tau,h})\mathds{1}_{A}]$ $=\mathbb{E}[\mathds{1}_{B}(Y_{\tau_{\infty},h})\mathds{1}_{A}]$
@@
where the last equality holds by the dominated convergence theorem, if the limit exists. To establish that the limit exists and compute it, we apply the usual de Finetti theorem – specificaly point $(i)$ of remark 12.27 in Klenke, (2008) with $f$ set to the identity map. This gives us that $\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\mathds{1}_{B}(Y_{\tau_{\infty},j})=\mathbb{E}[\mathds{1}_{B}(Y_{\tau_{\infty},h})\mid\Xi_{a_{h}}]=\Xi_{a_{h}}(B)$, where $\Xi_{a_{h}}(B)$ is the random variable $\omega\mapsto\Xi_{a_{h}}(\omega)(B)$. This in turn satisfies $\Xi_{a_{h}}(\omega)(B)=F(a_{h})(\omega)(B)=\kappa_{a_{h}}(\omega,B)$. This establishes that for any $F$-measurable set $A$ and Borel set $B$,
$\int\kappa_{F}(\omega,B)\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)=\int\kappa_{a_{h}}(\omega,B)\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)$
@@
    $\mathbb{E}[\mathds{1}_{B}(Y_{\tau,h})\mathds{1}_{A}]=\int\kappa_{F}(\omega,B)\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)=\int\kappa_{a_{h}}(\omega,B)\mathds{1}_{A}(\omega)d\mathbb{P}(\omega)$
@@ B.5 Establishing conditional independence of rewards
We will show this by induction on $H$. This clearly holds for $H=1$ by section B.3 above. Now assume that this holds for H=l1H=l-1.
@@
Showing that the empirical average is the conditional expectation: Again, consider a different infinite trajectory $\tau_{l}=(a_{l},a_{l},\dots)$. Again, by coherence, $Y_{\tau_{l},l+j}=Y_{\tau^{\prime},l+j}$ for all $j\geq 0$. From the usual de Finetti theorem, specifically point (i) of remark 12.27 in Klenke, (2008), we have that $\frac{1}{m}\sum_{j=1}^{m}f_{l}(Y_{\tau_{l},j})\to\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]$. We can then observe that by general properties of convergence, the following holds.
@@
Since both $U$ and $\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]$ are now $F$ measurable, we can apply the induction hypothesis to $\tau^{\prime}$ truncated at l1l-1 and conclude that
@@ B.6 A de Finetti theorem for TACDPs
We consider contexts and define contextual decision processes in a manner agnostic to context transitions. That is, the process only carries the data of how rewards are generated from given sequences of contexts and actions, while the context transitions themselves may be generated by a different process.
Definition 6.
It is natural to isolate away context dynamics for processes like stochastic contextual bandits, in which actions do not affect the context transitions – in contrast to transition-aware definitions in Jiang et al., (2016).
@@
A latent contextual bandit is a stateless decision process equipped with a random measure-valued function $F:\Omega\to((\mathcal{X}\times\mathcal{A})\to\mathcal{P}(\mathbb{R}))$ so that for any $H$ and context-action sequence $((x_{1},a_{1}),\dots(x_{H},a_{H}))$, the rewards $(Y_{1},\dots Y_{H}):=\mathcal{F}_{H}((x_{1},a_{1}),\dots(x_{H},a_{H}))$ are independent conditioned on $F$. Moreover, the conditional distribution $\mathcal{L}[Y_{h}\mid F]=F((x_{h},a_{h}))$ for all $h\leq H$. [^8] We abuse notation twice here. First, we write $F((x_{h},a_{h})):=(\omega\mapsto F(\omega)((x_{h},a_{h})))$. Second, as the regular conditional distribution $\mathcal{L}[Y_{h}\mid F]$ is a kernel that maps from $\Omega\times\mathcal{B}\to\mathbb{R}$, we view $F((x_{h},a_{h}))$ as its curried map $(\omega,B)\mapsto F((x_{h},a_{h}))(\omega)(B)$. A discussion of issues like measurability and well-definedness is in Appendix B.2.
@@ Appendix C Proofs for SOLD
With probability at least $1-\delta/2$, we have that $\|\overline{\textbf{M}}_{N}-\mathbb{E}[\textbf{M}_{1}]\|_{2}\leq\Delta_{M}$ with
@@
Now since $\|\textbf{M}_{n}^{2}\|_{2}\leq\|\textbf{M}_{n}\|_{2}^{2}\leq R^{4}\left(2+\frac{H}{4\mu}\right)^{2}$, we have that $\|\textbf{M}_{n}^{2}-\mathbb{E}[\textbf{M}_{1}^{2}]\|_{2}\leq 2R^{4}\left(2+\frac{H}{4\mu}\right)^{2}$ by the matrix Hoeffding bound, we have that with probability $1-\delta/2$,
@@
Further, by the matrix Bernstein inequality, we have that with probability $1-\delta/2$,
@@
Combining the two results and using a union bound, we get that with probability $1-\delta/2$
@@
With probability $1-\delta/4$, for $i=1,2$, we have that $\|\overline{\textbf{D}}_{N,i}-\mathbb{E}[\textbf{D}_{n,i}]\|_{2}=\left\|\frac{1}{N}\sum_{n=1}^{N}\textbf{D}_{n,i}-\mathbb{E}[\textbf{D}_{n,i}]\right\|_{2}\leq\Delta_{D}$ with
@@
Now note that by inequality (1.1) from Wei et al., (2005), we have that
@@
First note that $\mathbb{E}[\textbf{D}_{N,1}]^{-1}\mathbb{E}[\textbf{M}_{N,1}]\mathbb{E}[\textbf{D}_{N,2}]^{-1}=\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. Also recall that $\hat{\textbf{U}}$ is given by the top-$d_{K}$ eigenvectors for $\mathbb{E}[\textbf{D}_{N,1}^{-1}\textbf{M}_{N,1}\textbf{D}_{N,2}^{-1}]$. This means that there is a $d_{K}\times d_{K}$ unitary matrix W such that $\textbf{U}_{\star}\textbf{W}$ forms the eigenvectors for $\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. This means that by the Davis-Kahan theorem for statisticians Yu et al., (2014) and Lemma 5, we have that with probability $1-\delta$
@@
Bounding $\Delta_{\operatorname{off}}$: Now note that for large enough $N$, $\|\mathbb{E}[\textbf{D}_{n,1}]^{-1}\|_{2}\Delta_{D}=\sqrt{\frac{8\log(d_{A}/\delta)}{\lambda_{A}N}}\leq\frac{1}{2}$. In particular, by applying inequality (1.1) from Wei et al., (2005), we have that $B_{D}\leq 2\|\mathbb{E}[\textbf{D}_{n,1}]^{-1}\|_{2}=\frac{2}{\lambda_{A}}$. This already gives us the much simpler expression
@@ Appendix D Proofs for LOCAL-UCB
  • Note that with probability $1-\delta/3$, $\textbf{U}_{\star}$ satisfies $\|\hat{\textbf{U}}^{\top}\textbf{U}_{\star}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}$.
@@
$\stackrel{{\scriptstyle(i)}}{{=}}\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\tilde{\textbf{U}}_{t}\tilde{\textbf{U}}_{t}^{\top}\tilde{\bm{\beta}}_{t}-\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\bm{\beta}_{\star}$
@@
$\stackrel{{\scriptstyle(ii)}}{{=}}RT\Delta_{\operatorname{off}}+\sum_{t=1}^{T}\alpha_{2,t}\|\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\|_{\textbf{V}_{1,t}^{-1}}$
@@
$\stackrel{{\scriptstyle(iv)}}{{=}}O\left(Rd_{K}\sqrt{T\log\left(T/\delta\right)}+\frac{R\lambda_{D}^{3}}{\lambda_{\min}}\sqrt{\frac{T^{2}d_{K}d_{A}\log(d_{A}/\delta)}{N}}\right)$